OpenStudy (anonymous):

Solve for x: 2 over 5x + 2 = 10

jimthompson5910 (jim_thompson5910):

Is this \[\large \frac{2}{5}x + 2 = 10\] or is this \[\large \frac{2}{5x} + 2 = 10\] or is this \[\large \frac{2}{5x+2} = 10\] ???

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

It is equal to -9/25

jimthompson5910 (jim_thompson5910):

but which one is it?

OpenStudy (anonymous):

is it x=8

jimthompson5910 (jim_thompson5910):

no I meant the original problem

OpenStudy (anonymous):

x=-9/25.

jimthompson5910 (jim_thompson5910):

draw it out for us